OpenStudy (anonymous):

Solve each equation. Check each solution. 1. 2/( x + 3 ) + 5/( 3 - x ) = 6/( x^2 - 9 ) I got x= 9/7. When I plugged in that value for x, it doesn't equal each other so it has no solution, unless I did the problem wrong.

OpenStudy (anonymous):

\[\frac{ 2 }{ x+3 } + \frac{ 5 }{ 3-x } = \frac{ 6 }{ x ^{2} - 9 }\]

OpenStudy (anonymous):

first do the left side

OpenStudy (anonymous):

{2(3-x) +5(3+x)}/(9-x^2) =-6/(9-x^) 6-2x+15+5x=-6 3x+21=-6 3x=-27 x=-9

OpenStudy (anonymous):

I see what I did wrong. Thank you for your help!